# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

### v1

@@ 3.2. Results
The impact of genetic parameters like the secondary feature heritabilities and focal trait communality is clear (Figure 2). With the exception of the low focal trait communality settings, the CV1 accuracy of most multivariate approaches is, at worst, on par with that of the univariate model, across the range of focal trait heritabilities. Furthermore, CV1 accuracies of all methods, except the neural networks, converge to more or less the same point ($\rho\approx 0.65$) as $h^{2}(y)$ increases, highlighting that secondary features bring the biggest benefit when the focal trait is hard to measure accurately.
Most methods significantly outperform the univariate model in CV2, provided the quality of the secondary features is high enough. Looking at the setting with the highest quality secondary features, i.e., $h^{2}(s)=0.9$ and $\bm{\lambda}_{y}\bm{\lambda}_{y}^{\top}=0.8$, improvements with respect to the univariate approach reach approximately $40\%$ to $90\%$ depending on the focal trait heritability.
Regarding the secondary feature quality, the extent to which $h^{2}(s)$ and the focal trait communality impact performance differs between methods and also seems to depend on the focal trait heritability. Consider the cases where $h^{2}(y)=0.1$, for example. In this setting the heritability of secondary features has a negligible impact on the performance of gfBLUP and MegaLMM. In the case of lsBLUP, $h^{2}(s)$ has a small impact in CV1 when the focal trait communality is high, while it is relevant in combination with all communalities in CV2. Furthermore, significant overfitting seems to occur for lsBLUP in CV2 when both $\bm{\lambda}_{y}\bm{\lambda}_{y}^{\top}$ and $h^{2}(s)$ are low. For low focal trait heritabilities, siBLUP accuracy paradoxically decreases with increasing secondary feature heritabilities, in both CV1 and CV2. This confusing behaviour of siBLUP disappears when moving to a more realistic $h^{2}(y)=0.5$. The neural network approach performs significantly worse in both CV1 and CV2, next to having extremely inconsistent performance. Overall, both MegaLMM and gfBLUP seem to capture the factor structure underlying the data to the highest degree. Both approaches perform almost identically when averaging all results. The other approaches clearly react differently to complex interactions between genetic parameters like heritabilities and communalities.
4. Analysis of CIMMYT wheat data
In addition to the extensive simulations provided in the previous section we also evaluated the different approaches in terms of accuracy and interpretability using CIMMYT wheat data [5]. The following two sections provide details on the dataset, pre-processing, and results.
4.1. Data description and pre-processing
The CIMMYT wheat dataset contains measurements on grain yield and crop canopy hyperspectral reflectivity at $62$ wavelengths between $398$nm and $847$nm. The data was collected for four breeding cycles and five managed treatments within each cycle. We limited ourselves to a subset of the data to compare gfBLUP and the other methods. The subset we analyzed (breeding cycle 2014–2015, treatment B5IR) contains $39$ trials, $1094$ unique genotypes, and reflectance measurements at ten timepoints for a total of $620$ timepoint-wavelength combinations (referred to as the secondary features). The ten timepoints span the vegetative, heading (flowering), and grain-filling growth phases. Each trial had $28$ trial-specific genotypes and two check genotypes that occurred in every trial, all replicated three times. Trials were structured according to a resolvable incomplete block design. Marker data was available for $1033$ of the $1094$ genotypes. One genotype had missing yield, leaving us with $1032$ genotypes for the analyses. We first resolved the experimental design for each secondary feature using the following linear mixed model:
(12) $y_{(i|j)kl}=\mu+g_{j}+t_{k}+r_{i(k)}+b_{l(ik)}+\epsilon_{(i|j)kl}\text{,}$
where $y_{(i|j)kl}$ is yield or any of the $620$ secondary features for replicate $i$ given genotype $j$ in trial $k$, and block $l$. The overall mean is given by $\mu$ and $g_{j}$ is the fixed genotypic effect. The random trial effect $t_{k}$ is assumed to be independently, identically, and normally distributed with mean zero and variance $\sigma_{t}^{2}$: $t_{k}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Similarly, $r_{i(k)}\sim\mathcal{N}(0,\sigma_{r}^{2})$, $b_{l(ik)}\sim\mathcal{N}(0,\sigma_{b}^{2})$, and $\epsilon_{(i|j)kl}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ for the random replicate and block effects, as well as the residual effect. We then created the pseudo-CRD data by simply adding the estimated genetic effects and residuals from model (12):
$y_{q|j}=\mu+\hat{g}_{j}+\hat{\epsilon}_{(i|j)kl}\text{.}$
Note that due to the elimination of the design factors we can replace the indices $i$, $k$, and $l$ by a single index $q$ that denotes replicate $q$ of genotype $j$. We created $250$ datasets by randomly sampling $26$ trials as the training set and the remaining $13$ trials as the test set. We scaled and centered yield as well as the hyperspectral features separately for the training and test set prior to analyzing the data. We calculated the prediction accuracy as the correlation between predictions and the BLUEs ($\hat{g}_{j}$) obtained from model (12). We used a redundancy filtering threshold $\tau=0.95$ for all multivariate approaches as this provides a good trade-off between excluding redundant features and maintaining enough features to identify latent factors. Details on the methods other than gfBLUP can be found in SM Section S.5.
4.2. Results
Overall, all multivariate analyses except for deep learning performed practically the same with the CV2 predictions being more accurate than the CV1 predictions (Figure 3). None of the methods outperformed the univariate gBLUP model in CV1. Deep learning performed significantly worse than the other multivariate approaches, likely due to the relatively small dataset. Section S.8 of the SM provides results for a more practical scenario where only test set data measured during the vegetative phase is included in the CV2 predictions. Accuracies in this scenario are lower, but multivariate methods still outperform the univariate model showing that integration of genomic and phenomic data is promising even in more restrictive, realistic settings. We retained between $155$ and $172$ features after filtering using a threshold of $\tau=0.95$. The pattern of wavelengths that were retained was relatively constant across the ten dates, with most retained wavelengths representing the blue part of the spectrum. Reflectivities around 525nm, 700nm, and at 847nm are typically retained as well. A graphical overview can be found in Section S.9 of the SM.
@@ 4.2. Results
Figure 3. Prediction accuracies for several methods in the CV1 and CV2 scenarios for the wheat yield and hyperspectral reflectivities dataset. Redundancy filtering was carried out for all multivariate analyses with a threshold $\tau=0.95$.
@@
Figure 4. Loadings for the $62$ wavelengths and three factors found using gfBLUP and the hyperspectral data from 10-03-2015. Estimated genetic correlations between the three factors and yield are also shown.
@@
We reanalyzed the data from one of the timepoints (10-03-2015) to assess the interpretability of gfBLUP and the other approaches. We omitted the redundancy filtering step, resulting in $62$ wavelengths. Figure 4 shows the gfBLUP factor loadings obtained through this analysis. All three factors were highly heritable at $H^{2}=0.79$ for factor $1$ and $H^{2}=0.72$ for factor $2$, and $H^{2}=0.55$ for factor $3$, where $H^{2}$ denotes the wide-sense heritability as variances were estimated using sums of squares and are thus not additive only. Genetic correlations between the factors and yield were $-0.63$ for factor $1$, $0.27$ for factor $2$, and $0.15$ for factor $3$. Through these correlations and the factor loadings, it is immediately clear that higher yielding genotypes tended to have lower reflectivities in the green-red part of the spectrum and higher reflectivities in the near-infrared (NIR) part. These results are reminiscent of the normalized difference vegetation index (NDVI) that is highly correlated with yield and is high itself for crop canopies with high NIR reflectivity and low red light reflectivity [35]. We provide details on the interpretability of the other approaches in SM Section S.6. We did not include deep learning here as it was impossible to extract any biological meaning from the parameters of the neural networks. We provide some data on the computational complexity of gfBLUP and other methods in SM Section S.7.
@@ 5. Discussion
With the recent innovations in HTP techniques and the corresponding increase in data available to plant breeders, the development of methodology for integrating this data has become increasingly relevant. We have shown that the proposed methodology gfBLUP performs on par or better than alternatives and is able to integrate secondary HTP data efficiently without loss of interpretability.
@@
We used simulations to evaluate the impact of several genetic parameters on model performance. Of these parameters, the focal trait heritability is of particular interest. One of the key motivations for the use of secondary data in multivariate genomic prediction is the belief that this data can improve predictions in cases where the focal trait phenotype has low heritability [11]. The proposed gfBLUP performs well in such cases due to its unsupervised nature. It does not suffer from overfitting at low focal trait heritabilities like some alternative approaches.
@@
We alse evaluated interpretability by analyzing a wheat yield and hyperspectral reflectivity dataset. We found that two of the three obtained factors were responsible for the (co)variation in reflectivities of visible and NIR light. Combined with the genetic correlations between these factors and yield we found a pattern similar to the NDVI which is known to be a good indicator of yield [36]. As large-scale data collection becomes increasingly accessible, the complexity of datasets and the importance of understanding biological implications increases as well [37]. Given the results obtained here, incorporating generative factor analysis in genomic prediction seems an effective way of tackling these challenges.
@@
There are several options to extend the proposed methodology. While the unsupervised approach of gfBLUP worked well here, there are benefits to using a supervised approach. One way of achieving this would be to use a factor regression approach [13]. Conditioning on the prediction target may lead to higher accuracy when the focal trait training data is of sufficient quality. Another option would be to consider several ways of integrating multiple data modalities as opposed to using a single (e.g., hyperspectral) modality as we did here. There are various methods that can generally be categorized into early integration, intermediate integration, or late integration [38]. These integration approaches each try to balance parsimony and accuracy. It may be fruitful to develop methodology that finds the right balance between the parsimony and accuracy of different integration approaches within the context of genomic prediction. Finally, there is the issue of multi-year, multi-environment trials [39]. By analyzing a trait of interest measured in several environments using a multivariate model, gfBLUP could reveal latent drivers of environment-specific variety performance. Combining this approach with dynamic factor analysis [40] would provide a promising framework to model the main source of data in large breeding programs.
@@
In any case, the gfBLUP methodology we propose here is a robust and relatively efficient approach to obtaining multivariate genomic predictions with several promising options regarding extensions and further research.
6. Data availability
@@
Scripts to generate the simulated and hyperspectral datasets, as well as the results presented in this paper, are available at https://github.com/KillianMelsen/gfBLUP_2024. The gfBLUP methodology is implemented in an R-package available at https://github.com/KillianMelsen/gfBLUP. The hyperspectral dataset is available upon reasonable request from J. Crossa.
@@
S.1. Notation
We use $\circ$ and $\otimes$ to denote the Hadamard and Kronecker products. We use $diag(s_{11},\dots,s_{pp})$ to refer to a $p$-dimensional diagonal matrix with elements $s_{11},\dots,s_{pp}$. We denote the column-wise vectorization of a matrix by $\operatorname{vec}(\mathbf{A})$ and use $\operatorname{vec}^{-1}(\mathbf{A})$ for the reverse operation. The cardinality of a set $\mathcal{A}$ is denoted by $\text{card}(\mathcal{A})$. Some further information and examples on the partitioning of data and covariance matrices, as well as some general background information on genomic prediction, can be found in SM Section S.2.
S.2. General genomic prediction setting
We more or less follow the notation of Runcie and Cheng [1] and Arouisse et al. [2]. Let $\mathbf{Y}\in\mathbb{R}^{n\times(p+1)}$ be the data matrix containing unreplicated, mean-centered, and scaled measurements on $p$ secondary features and the focal trait for $n$ genotypes. We can divide the $n$ genotypes into a test and training set so that $n=n_{t}+n_{o}$ ($t$ for test and $o$ for the observed training set). The matrix $\mathbf{Y}$ can then be partitioned as:
@@
where $\mathbf{Y}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}$ is vectorized. Expression (S2) can be evaluated using a fast algorithm that decreases the time complexity for the calculation of the focal trait training set BLUPs $\hat{\mathbf{g}}^{(CV1)}_{fo}$ from $O(n_{o}^{3}p^{3})$ to $O(n_{o}p^{2}+p^{3})$ by avoiding the naive inversion of $\hat{\mathbf{V}}$ [3].
The CV2 multivariate predictions can be obtained directly [2] or through a two-step approach [1]. We prefer the two-step approach because the fast algorithm can be used to obtain all CV1 training set BLUPs, $\hat{\mathbf{G}}^{(CV1)}_{.o}$:
@@
where $\mathbf{g}_{s(j)}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{ss}^{g})$ and $\bm{\epsilon}_{s(i|j)}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{ss}^{\epsilon})$. All $r$ residual vectors for the replicates of genotype $j$ are assumed iid, meaning $\text{var}(\bm{\epsilon}_{s(i|j)})=\text{var}(\bm{\epsilon}_{s(i^{\prime}|j)})=\bm{\Sigma}_{ss}^{\epsilon}$ and the cross-covariance matrix for any two replicates of genotype $j$, $\text{cov}(\bm{\epsilon}_{s(i|j)},\bm{\epsilon}_{s(i^{\prime}|j)})$, is a zero matrix. As a result,
@@
S.4. Efficient penalty optimization
We use Brent’s algorithm [11] implemented in the optimize() function in R to find the optimal values for $\vartheta_{g}$ and $\vartheta_{\epsilon}$ as described in the main text. In a naive implementation we obtain the mean cross-validated negative log-likelihood at each iteration as:
@@
S.5. Alternative methods
S.5.1. Univariate gBLUP
@@
where $\mathbf{y}$ is the $n\times 1$ vector containing the mean-centered and scaled focal trait phenotypes, $\mathbf{Z}$ is the incidence matrix linking individuals to the BLUPs in $\mathbf{u}\sim\mathcal{N}(\mathbf{0},\sigma_{g}^{2}\mathbf{K})$, and $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ contains the residuals. We used the R-package rrBLUP version 4.6.1 [4] to fit univariate gBLUP models (see S1). The marker-based kinship matrix was created using the R-package statgenGWAS version 1.0.9.
S.5.2. siBLUP
The first alternative multivariate approach that we included is a modified version of the bivariate gBLUP model using a regularized selection index as proposed by Lopez-Cruz et al. [5] and also implemented by Arouisse et al. [2]. In short, siBLUP uses a regularized phenotypic covariance matrix of the secondary features and their genetic covariances with the focal trait to obtain a weight for every secondary feature. We use these weights to calculate a single linear combination of the secondary features for each individual termed the (regularized) selection index. We then use this index as the single secondary feature in the multivariate gBLUP equations (SM Section S.2).
Like gfBLUP, the modified version of siBLUP takes data containing replicates from a completely randomized design. Next, we obtain $\hat{\bm{\Sigma}}_{ss}^{g}$ and $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ using sums of squares, after which we redundancy filter the dataset so that no two of the remaining secondary features have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding a threshold $\tau$.
@@
S.5.3. lsBLUP
lsBLUP also reduces the secondary features to a single dimension [2]. For lsBLUP, this secondary feature is a LASSO regression prediction of the focal trait using the redundancy filtered set of secondary features.
We use sums of squares obtained from replicated completely randomized design data to estimate $\bm{\Sigma}_{ss}^{g}$. We then use this genetic covariance matrix of the original secondary features (converted to the correlation scale) to redundancy filter the data. After filtering, the LASSO regression coefficients $\hat{\bm{\beta}}^{LS}$ are obtained using the R-package glmnet version 4.1.6 [6]. To do so, we first reduce the redundancy filtered data to genotypic means. We then use the cv.glmnet() function to determine the optimal (i.e., minimum mean cross-validated error) LASSO penalty $\lambda^{\ddagger}$ using 5-fold cross-validation, thereby obtaining the vector of coefficients given this optimal penalty $\hat{\bm{\beta}}^{LS}(\lambda^{\ddagger})$.
If all coefficients have been regularized to 0, lsBLUP returns simple univariate gBLUP predictions for the test set using the rrBLUP R-package [4]. If not, we calculate plot-level LASSO predictions:
@@
and compute the CV1 or CV2 test set focal trait BLUPs as described for siBLUP and gfBLUP, using $\hat{\mathbf{y}}_{LSP}$ as the single secondary feature.
S.5.4. MegaLMM
MegaLMM was introduced as a method to make genomic predictions using many secondary features [7]. In essence, it decomposes the $p$ original features into two parts. One part is projected onto $M$ latent factors while the other part represents the $p$ independent errors for each secondary feature. Independent linear mixed models are then fit for these $M+p$ decorrelated traits. For the full details on MegaLMM see Runcie et al. [7].
To be able to make consistent comparisons, we redundancy filtered the full set of secondary features as we did for other methods prior to calculating the BLUEs and running MegaLMM. We initialized and ran MegaLMM with default settings except for the following parameters: for the MegaLMM_control() function drop0_tol was set to $10^{-10}$, h2_divisions was set to $20$, burn was set to 00 (we did not use the built-in function to automatically discard burn-in iterations), and thin was set to $2$. The number of modeled latent factors $M$ (argument K) was set to $16$ for the simulated data and $20$ for the hyperspectral data. For the priors we again mostly used the default settings with two exceptions. For tot_Y_var, we set nu to $10$ for both simulated and hyperspectral data. We set nu for tot_F_var to $100{,}000$ and $20$ for the simulated and hyperspectral data, respectively.
After initializing MegaLMM we first ran $1000$ burn-in iterations, reordering the factors after each $100$ iterations, for both simulated and hyperspectral data. We then ran another 2000 iterations for the simulated data and 500 iterations for the hyperspectral data. Posterior predictions for each genotype were calculated at each iteration, with the final predictions being equal to the posterior means. We used version 0.1.0 of the MegaLMM R-package [8].
S.5.5. Deep learning

### v3

@@ 4.2. Results
The relative performance of the different models clearly depends on the subset of the data and thus the managed treatment (Figure 3). Our proposed pipeline performs best in the HEAT treatment while lsBLUP performs best for B5IR. All multivariate approaches show CV1 performances that are more or less similar to the performance of the univariate model, with the exception of the deep learner which performs significantly worse in all scenarios. For the other approaches CV2 performance is clearly higher than the univariate model, especially in the B5IR treatment. Only including test set hyperspectral measurements from the vegetative phase provided a benefit in B5IR but not in the HEAT treatment, possibly due to the fact that there was only a single measuring point during the vegetative phase for HEAT.
At this point we would like to point out that by using a two-stage modeling approach, i.e., performing the spatial correction first and then using the BLUEs from the first stage in the univariate model, we are likely slightly underestimating the univariate model’s accuracy. It would be trivial to fit a single-stage univariate model in a real world application where a complete trial would be the training set. This single stage model would perform spatial correction and genomic prediction for test set genotypes not present in the trial simultaneously. Unfortunately, in a cross-validation setting where training and test set are obtained from the same trial, this is infeasible.
The pattern of wavelengths that were retained after redundancy filtering was not constant across the ten dates for B5IR and four dates for HEAT, although most retained wavelengths represented the blue, red, and near infra-red part of the spectrum. A graphical overview of what features were retained after filtering can be found in SM Section 7.
@@
Figure 3. Prediction accuracies for several methods in the CV1, CV2, and CV2VEG scenarios for two managed treatments for the wheat yield and hyperspectral reflectivities dataset. Redundancy filtering was carried out for all multivariate analyses with a threshold $\tau=0.95$.
@@
Figure 4. Loadings for the $62$ wavelengths and three factors found using glfBLUP and the hyperspectral data from 10-03-2015. Estimated genetic correlations between the three factors and yield are also shown.
@@
We reanalyzed the data from one of the timepoints (10-03-2015) of the B5IR treatment to assess the interpretability of glfBLUP and the other approaches. We omitted the redundancy filtering step, resulting in $62$ wavelengths. Figure 4 shows the glfBLUP factor loadings obtained through this analysis. All three factors were highly heritable at $H^{2}=0.98$ for factor $1$, $H^{2}=0.96$ for factor $2$, and $H^{2}=0.95$ for factor $3$, where $H^{2}$ denotes the wide-sense heritability as variances were estimated using sums of squares and are thus not additive only. Genetic correlations between the factors and yield were $-0.67$ for factor $1$, $0.26$ for factor $2$, and $0.16$ for factor $3$. Through these correlations and the factor loadings, it is immediately clear that higher yielding genotypes tended to have lower reflectivities in the green-red part of the spectrum or higher reflectivities in the near-infrared (NIR) part. These results are somewhat reminiscent of the normalized difference vegetation index (NDVI) that is highly correlated with yield and is high itself for crop canopies with high NIR reflectivity and low red light reflectivity (Teal et al., 2006). This showcases the ability of glfBLUP to reveal biologically relevant patterns in high-throughput phenotyping data. We provide details on the interpretability of the other approaches in SM Section 5. We did not include deep learning here as it was impossible to extract any biological meaning from the parameters of the neural networks. We provide some data on the computational complexity of glfBLUP and other methods in SM Section 6.
@@ 5. Discussion
With the recent innovations in HTP techniques and the corresponding increase in data available to plant breeders, the development of methodology for integrating these data has become increasingly relevant. We have developed a pipeline approach that combines classical factor analysis, which is typically applied to phenotypic data, with methods for high-dimensional settings and elements from mixed modeling approaches for genomic prediction. We have also adapted several of these elements to facilitate the modeling of genetic and non-genetic sources of variation, showing the strength of combining existing methodologies and making critical changes that allow these methods to be applied to different types of data. We have shown that the proposed methodology glfBLUP performs on par or better than alternatives and is able to integrate secondary HTP data efficiently without loss of interpretability.
@@
We used simulations to evaluate the impact of several genetic parameters on model performance. Of these parameters, the focal trait heritability is of particular interest. One of the key motivations for the use of secondary data in multivariate genomic prediction is the belief that these data can improve predictions in cases where the focal trait phenotype has low heritability (Lopez-Cruz et al., 2020). The proposed glfBLUP performs well in such cases due to its unsupervised nature. It does not suffer from overfitting at low focal trait heritabilities like some alternative approaches seem to do.
Our approach also does not rely as heavily on hyperparameter tuning as approaches such as deep learning. The performance of the neural networks on both the simulated data as well as the real data was significantly lower than any of the other models. This could very well be a result of the chosen architecture and hyperparameters not being optimal. This result immediately showcases the difficulty with many machine learning approaches. Proper hyperparameter tuning is either time consuming, computationaly expensive, or both. We thus view the minimal required tuning in glfBLUP as a major advantage.
@@
Another advantage of glfBLUP over machine learning approaches is the interpretability. We evaluated interpretability by analyzing a wheat yield and hyperspectral reflectivity dataset. We found that two of the three obtained factors were responsible for the (co)variation in reflectivities of visible and NIR light. Combined with the genetic correlations between these factors and yield we found a pattern similar to the NDVI which is known to be a good indicator of yield (Mkhabela et al., 2011). As large-scale data collection becomes increasingly accessible, the complexity of datasets and the importance of understanding biological implications increases as well (Danilevicz et al., 2022). Given the results obtained here, incorporating generative factor analysis in genomic prediction seems an effective way of tackling these challenges.
Regarding factor analysis, we opted to use maximum likelihood estimation and a modified Thomson regression approach to obtain estimates of the loadings, unique variances, and factor scores (Jöreskog, 1967; Thomson, 1939). Other approaches based on principal component analysis exist as well (e.g., Zhang et al., 2022; Fan et al., 2021) and may be an alternative to obtain estimates in cases where the data contains heteroskedastic noise.
There are other implementations of the factor model too (Smith et al., 2001; Tolhurst et al., 2019). These linear mixed models provide an attractive single-step approach to modelling multi-trait data by assuming a factor analytic structure for the genetic covariance matrix of traits and estimating any parameters using restricted maximum likelihood. Such single-stage approaches have a number of advantages over using a two-stage approach and pseudo-CRD data. Correcting data for nuisance effects such as spatial trends can be done most effectively by using estimates obtained from a full model including all effects (Piepho et al., 2013). Furthermore, a single-stage approach will result in valid standard errors for all predictions. While the single-step nature of linear mixed models is thus attractive, their computational complexity means they are typically not used for data with hundreds or potentially thousands of traits (Runcie et al., 2021). We demonstrated that using estimates from reduced models to produce pseudo-CRD data for further modeling works well for prediction. The approach we present here combines the attractive properties of factor analysis with a pipeline for prediction that avoids the computationally complex characteristics of the single-step mixed model approaches needed to obtain quantities such as standard errors.
@@
We recognize several options to extend the proposed methodology. Our motivation for using an unsupervised dimensionality reduction approach was based on the idea that secondary phenotypes are often collected when the focal trait has a low heritability (Sun et al., 2017). In these cases, the noisy focal trait data may negatively impact the estimation of latent factors. While the unsupervised approach of glfBLUP worked well here, likely due to the noisy nature of the focal trait data, there are benefits to using supervised approaches like the other models we evaluated. We could include the focal trait in the factor analysis we used to reduce the dimensionality, much like MegaLMM. The factor analysis step could also be replaced by other supervised dimensionality reduction methods such as partial least squares regression to automatically obtain a small set of features that are maximally correlated with the focal trait (Colombani et al., 2012). Another way of achieving this would be to model secondary features as predictors instead of multivariate response variables by using a factor regression approach (Münch et al., 2022). Conditioning on the prediction target may lead to higher accuracy when the focal trait training data is of sufficient quality. Another option would be to consider several ways of integrating multiple data modalities as opposed to using a single (e.g., hyperspectral) modality as we did here. There are various methods that can generally be categorized into early integration, intermediate integration, or late integration (Pavlidis et al., 2001). These integration approaches each try to balance parsimony and accuracy. It may be fruitful to develop methodology that finds the right balance between the parsimony and accuracy of different integration approaches within the context of genomic prediction. Finally, there is the issue of multi-year, multi-environment trials (Cullis et al., 2014). By analyzing a trait of interest measured in several environments using a multivariate model, glfBLUP could reveal latent drivers of environment-specific variety performance. Combining this approach with dynamic factor analysis (Molenaar, 1985) would provide a promising framework to model the main source of data in large breeding programs.
@@
In any case, the glfBLUP methodology we propose here is a robust and relatively efficient approach to obtaining multivariate genomic predictions with several promising options regarding extensions and further research.
Author Contributions
Conceptualization: C.F.W.P. and W.K.; methodology and software: K.A.C.M. and J.F.K.; writing–original draft preparation: K.A.C.M.; writing–review and editing: K.A.C.M., J.F.K., J.C., M.R.K., F.A.E., W.K., C.F.W.P. All authors have read and agreed to the published version of the manuscript.
Conflict of Interest
The authors have declared no conflict of interest.
Data Availability Statement
@@
Code to generate the data and results presented in this paper are available at https://github.com/KillianMelsen/glfBLUP_2025. The full hyperspectral dataset is available upon reasonable request from J. Crossa. The proposed approach glfBLUP has been implemented in an R-package available at https://github.com/KillianMelsen/glfBLUP.
Appendix: Additional simulation results
Here we provide some additional results using the simulated data (Figure 5). Compared to the $h^{2}(s)=0.7$ setting the conclusions for $h^{2}(s)=0.5$ and $h^{2}(s)=0.9$ are largely similar. Overall, there is a slight decrease in accuracy when the heritability of the secondary features drops, but the focal trait unique variance and thus the degree to which the focal trait is genetically correlated to the secondary features is far more important.
Figure 6 shows results obtained when simulating data using a low-rank structure for the residual correlations instead of small, random correlations (see Figure 7 and Section 3.1 for details on the simulation process).
Figure 5. Prediction accuracies for several methods in the CV1 and CV2 scenarios as we vary the focal trait communality, focal trait heritability, and secondary feature heritabilities (random residual correlations). The benchmark predictions use the true simulated factor scores and covariance matrices for multivariate genomic prediction. Further details regarding these simulations can be found in Section 3.1.
Figure 6. Prediction accuracies for several methods in the CV1 and CV2 scenarios as we vary the focal trait communality, focal trait heritability, and secondary feature heritabilities (low-rank model for the residual correlations). The benchmark predictions use the true simulated factor scores and covariance matrices for multivariate genomic prediction. Further details regarding these simulations can be found in Section 3.1.
[…]
@@
S.1. Genomic prediction
We more or less follow the notation of Runcie and Cheng (2019) and Arouisse et al. (2021). Let $\mathbf{Y}\in\mathbb{R}^{n\times(p+1)}$ be the data matrix containing unreplicated, mean-centered, and scaled measurements on $p$ secondary features and the focal trait for $n$ genotypes. We can divide the $n$ genotypes into a test and training set so that $n=n_{t}+n_{o}$ ($t$ for test and $o$ for the observed training set). The matrix $\mathbf{Y}$ can then be partitioned as:
@@
where $\mathbf{Y}_{.o}-\mathbf{X}_{o}\hat{\bm{\beta}}$ is vectorized. Expression (S2) can be evaluated using a fast algorithm that decreases the time complexity for the calculation of the focal trait training set BLUPs $\hat{\mathbf{g}}^{(CV1)}_{fo}$ from $O(n_{o}^{3}p^{3})$ to $O(n_{o}p^{2}+p^{3})$ by avoiding the naive inversion of $\hat{\mathbf{V}}$ (Dahl et al., 2013).
The CV2 multivariate predictions can be obtained directly (Arouisse et al., 2021) or through a two-step approach (Runcie and Cheng, 2019). We prefer the two-step approach because the fast algorithm can be used to obtain all CV1 training set BLUPs, $\hat{\mathbf{G}}^{(CV1)}_{.o}$:
@@
where $\mathbf{g}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{ss}^{g})$ and $\bm{\epsilon}_{s(i|j)}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{ss}^{\epsilon})$. All $r$ residual vectors for the replicates of genotype $j$ are assumed iid, meaning $\text{var}(\bm{\epsilon}_{s(i|j)})=\text{var}(\bm{\epsilon}_{s(i^{\prime}|j)})=\bm{\Sigma}_{ss}^{\epsilon}$ and the cross-covariance matrix for any two replicates of genotype $j$, $\text{cov}(\bm{\epsilon}_{s(i|j)},\bm{\epsilon}_{s(i^{\prime}|j)})$, is a zero matrix. As a result,
@@
S.3. Penalty optimization
We use Brent’s algorithm (Brent, 1973) implemented in the optimize() function in R to find the optimal values for $\vartheta_{g}$ and $\vartheta_{\epsilon}$ as described in the main text. In a naive implementation we obtain the mean cross-validated negative log-likelihood at each iteration as:
@@
S.4. Alternative methods
S.4.1. Univariate gBLUP
@@
where $\mathbf{y}$ is the $n\times 1$ vector containing the mean-centered and scaled focal trait phenotypes, $\mathbf{Z}$ is the incidence matrix linking individuals to the BLUPs in $\mathbf{u}\sim\mathcal{N}_{n_{g}}(\mathbf{0},\sigma_{g}^{2}\mathbf{K})$, and $\bm{\epsilon}\sim\mathcal{N}_{n}(\mathbf{0},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ contains the residuals. We used the R-package rrBLUP version 4.6.1 (Endelman, 2011) to fit univariate gBLUP models (see S1). The marker-based kinship matrix was created using the R-package statgenGWAS version 1.0.9.
S.4.2. siBLUP
The first alternative multivariate approach that we included is a modified version of the bivariate gBLUP model using a regularized selection index as proposed by Lopez-Cruz et al. (2020) and also implemented by Arouisse et al. (2021). In short, siBLUP uses a regularized phenotypic covariance matrix of the secondary features and their genetic covariances with the focal trait to obtain a weight for every secondary feature. We use these weights to calculate a single linear combination of the secondary features for each individual termed the (regularized) selection index. We then use this index as the single secondary feature in the multivariate gBLUP equations (SM Section S.1).
Like glfBLUP, the modified version of siBLUP takes data containing replicates from a completely randomized design. Next, we obtain $\hat{\bm{\Sigma}}_{ss}^{g}$ and $\hat{\bm{\Sigma}}_{ss}^{\epsilon}$ using sums of squares, after which we redundancy filter the dataset so that no two of the remaining secondary features have an absolute genetic correlation $|\rho^{g}|$ equal to, or exceeding a threshold $\tau$.
@@
S.4.3. lsBLUP
lsBLUP also reduces the secondary features to a single dimension (Arouisse et al., 2021). For lsBLUP, this secondary feature is a LASSO regression prediction of the focal trait using the redundancy filtered set of secondary features.
We use sums of squares obtained from replicated completely randomized design data to estimate $\bm{\Sigma}_{ss}^{g}$. We then use this genetic covariance matrix of the original secondary features (converted to the correlation scale) to redundancy filter the data. After filtering, the LASSO regression coefficients $\hat{\bm{\beta}}^{LS}$ are obtained using the R-package glmnet version 4.1.6 (Friedman et al., 2010). To do so, we first reduce the redundancy filtered data to genotypic means. We then use the cv.glmnet() function to determine the optimal (i.e., minimum mean cross-validated error) LASSO penalty $\lambda^{\ddagger}$ using 5-fold cross-validation, thereby obtaining the vector of coefficients given this optimal penalty $\hat{\bm{\beta}}^{LS}(\lambda^{\ddagger})$.
If all coefficients have been regularized to 0, lsBLUP returns simple univariate gBLUP predictions for the test set using the rrBLUP R-package (Endelman, 2011). If not, we calculate plot-level LASSO predictions:
@@
and compute the CV1 or CV2 test set focal trait BLUPs as described for siBLUP and glfBLUP, using $\hat{\mathbf{y}}_{LSP}$ as the single secondary feature.
S.4.4. MegaLMM
MegaLMM was introduced as a method to make genomic predictions using many secondary features (Runcie et al., 2021). In essence, it decomposes the $p$ original features into two parts. One part is projected onto $M$ latent factors while the other part represents the $p$ independent errors for each secondary feature. Independent linear mixed models are then fit for these $M+p$ decorrelated traits. For the full details on MegaLMM see Runcie et al. (2021).
To be able to make consistent comparisons, we redundancy filtered the full set of secondary features as we did for other methods prior to calculating the BLUEs and running MegaLMM. We initialized and ran MegaLMM with default settings. The number of modeled latent factors $M$ (argument K) was set to $16$ for the simulated data and $20$ for the hyperspectral data. For the priors we again used the default settings.
After initializing MegaLMM we first ran $1000$ burn-in iterations, reordering the factors after each $100$ iterations, for both simulated and hyperspectral data. We then ran another 500 iterations for the posterior samples. Posterior predictions for each genotype were calculated at each iteration, with the final predictions being equal to the posterior means. We used version 0.1.0 of the MegaLMM R-package (Runcie, 2022).
S.4.5. Deep learning